# Synergizing Motion and Appearance: Multi-Scale Compensatory Codebooks for Talking Head Video Generation

### v1

@@ A.2.2 More Experimental Results
Cross-identity Reenactment. Since there is no ground truth for cross-identity reenactment, we conduct a user study comparing our approach to recent state-of-the-art methods, including a GAN-based model (MRFA (Tao et al., 2024)) and a diffusion-based model (Follow-You-Emoji (FYE) (Ma et al., 2024)). We randomly selected 10 source-driving pairs and asked 30 participants to evaluate the generated videos based on appearance realism, motion naturalness, and overall quality. The results shown in Fig. 12 indicate that users prefer our method, confirming its superiority.
Table 3: Quantitative comparison with state-of-the-art methods for cross-identity reenactment on VoxCeleb1 dataset. AniPortrait (Wei et al., 2024) and Follow-Your-Emoji (FYE) (Ma et al., 2024) are trained on much larger-scale datasets and are not suitable for a direct comparison.
Method Training Dataset FID $\downarrow$ CSIM $\uparrow$ ARD $\downarrow$
AniPortrait (Wei et al., 2024) VFHQ (Xie et al., 2022), CelebV-HQ (Zhu et al., 2022) 66.61 0.7226 2.9146
FYE (Ma et al., 2024) HDTF (Zhang et al., 2021), VFHQ (Xie et al., 2022), their collected dataset (Ma et al., 2024) 60.05 0.7558 3.0822
FOMM (Siarohin et al., 2019b) VoxCeleb1 (Nagrani et al., 2017) 80.00 0.6010 1.8331
DaGAN (Hong et al., 2022) VoxCeleb1 (Nagrani et al., 2017) 85.32 0.5743 2.0604
MCNet (Hong & Xu, 2023) VoxCeleb1 (Nagrani et al., 2017) 82.72 0.5618 1.6970
MRFA (Tao et al., 2024) VoxCeleb1 (Nagrani et al., 2017) 77.63 0.5962 1.5903
Ours VoxCeleb1 (Nagrani et al., 2017) 76.47 0.6142 1.6234
@@
We also present a quantitative comparison with state-of-the-art methods for cross-identity reenactment in Tab. 3. We use FID (Heusel et al., 2017) for image quality evaluation, Average Rotation Distance (ARD) for motion transfer evaluation following Tao et al. (2024), and cosine similarity (CSIM) for identity preservation following Ha et al. (2020). Diffusion-based AniPortrait (Wei et al., 2024) and Follow-Your-Emoji (FYE) (Ma et al., 2024) are trained on much larger-scale datasets and are excluded from the comparison. Our method performs the best in terms of image quality and identity preservation, and is comparable in terms of motion transfer, confirming the effectiveness of our approach. Although MRFA (Tao et al., 2024) can transfer motion well, its output frame quality may be low and it may not preserve source identity effectively. Although diffusion-based methods can achieve even better image quality and identity preservation performance, they struggle to transfer motion faithfully, which leads to undesirable visual quality.
Table 4: Ablation study on the code allocation scheme. We present the results of different codebook splitting schemes.
@@
Effect of the code allocation scheme. We propose a novel code allocation scheme for motion and appearance codebooks that assigns different codes to corresponding scales. This allows certain codes to be shared across multiple scales, facilitating the transfer of information between them. To assess the effectiveness of our code allocation scheme, we conduct an ablation study, as shown in Tab. 4. We compare two alternative codebook splitting schemes: sharing all codes across all scales and splitting the codes equally among the scales. As demonstrated in Tab. 4, our code allocation scheme generally achieves the best overall performance. These results highlight the effectiveness and superiority of our proposed code allocation scheme.
A.3 Limitations
A limitation of our method is the identity leakage problem in cross-identity reenactment, where the face in the generated video tends to have a shape similar to that of the driving face rather than the source face. This issue arises from the keypoint-based motion flow estimator that we adopt to produce the initial coarse motion flow and the driving keypoints for multi-scale motion codebook compensation. Although this motion flow estimator is robust to non-facial motion, such as hair and neck movement, by learning unsupervised keypoints on talking heads, the keypoints also inherently model facial shapes, which leads to the entanglement of motion and shape. Relative motion transfer (Siarohin et al., 2019b) can alleviate this problem and is widely adopted by previous methods (Hong et al., 2022; Hong & Xu, 2023; Tao et al., 2024).

### v2

@@ 7.2.2 More Comparison Results
Figure 13: Qualitative comparison with more state-of-the-art approaches for (a) same-identity reconstruction and (b) cross-identity reenactment on VoxCeleb1 or examples from the corresponding papers or project pages for closed-source methods (i.e., OSFV [29], PECHead [9], and MegaPortraits [6]). Our method better mimics the driving motion and preserves more facial details.
@@
We also present quantitative comparison results for cross-identity reenactment in Tab. 3. We use FID [14] for image quality evaluation, Average Rotation Distance (ARD) for motion transfer evaluation following [25], and cosine similarity (CSIM) for identity preservation following [12]. Diffusion-based AniPortrait [32] and Follow-Your-Emoji (FYE) [19] are trained on much larger-scale datasets and are excluded from the comparison. Our method generally demonstrates the highest overall performance, confirming its effectiveness. LIA [31] is slightly better in image quality and identity preservation, as it uses latent codes instead of keypoints as the motion representation, which helps appearance preservation. However, its motion transfer quality is much worse. We also provide a qualitative comparison with LIA in Fig. 13 where our method mimics the driving motion better. Although MRFA [25] can transfer motion well, its output frame quality may be low and it may not preserve source identity effectively. Although diffusion-based methods [32, 19] can achieve even better image quality and identity preservation performance, they struggle to transfer motion faithfully, which leads to undesirable visual quality.
Comparison with more state-of-the-art approaches. We additionally provide comparisons with more state-of-the-art approaches, including an open-source method (i.e., LivePortrait [11]) and several closed-source methods (i.e., OSFV [29], PECHead [9], and MegaPortraits [6]). The qualitative comparison in Fig. 13 highlights the advantages of our method in the preservation of facial details and expression transfer. We also provide a quantitative comparison with the open-source LivePortrait [11] in Tab. 4. Although LivePortrait is trained on significantly larger datasets and is unsuitable for a direct fair comparison, our method can still outperform it on all metrics for same-identity reconstruction.
Method # Training Same-identity Reconstruction Cross-identity Reenactment
Video Frames FID $\downarrow$ PSNR $\uparrow$ $\mathcal{L}_{1}$downarrow LPIPS $\downarrow$ AKD $\downarrow$ AED $\downarrow$ FID $\downarrow$ CSIM $\uparrow$ ARD $\downarrow$
LivePortrait [11] 69M 48.11 22.94 0.0484 0.2213 1.5516 0.1602 75.95 0.7260 1.3497
Ours 4.3M 43.15 25.30 0.0355 0.1846 1.2039 0.1071 76.47 0.6142 1.6234
Table 4: Quantitative comparison with LivePortrait [11] on VoxCeleb1. LivePortrait, being trained on significantly larger data, is unsuitable for a direct comparison.
Inference speed. We evaluate the inference speed using an NVIDIA RTX 3090 and provide the results in Tab. 5. Our approach shows clear advantages upon recent state-of-the-art methods [25, 32, 19, 11], indicating its potential for real-time performance.
MRFA [25] AniPortrait [32] FYE [19] LivePortrait [11] Ours
FLOPs $\downarrow$ 403.05G 9.18T 15.04T 1.31T 352.91G
FPS $\uparrow$ 12.41 0.36 0.39 11.28 15.13
Table 5: Inference speed comparison.
7.2.3 Additional Ablation Study
Effect of the code allocation scheme. We propose a novel code allocation scheme for motion and appearance codebooks that assigns different codes to corresponding scales. This allows certain codes to be shared across multiple scales, facilitating the transfer of information between them. To assess the effect of our code allocation scheme, we conduct an ablation study and present results in Tab. 6. We compare with two alternative codebook splitting schemes: sharing all codes across all scales and splitting the codes equally among the scales. As demonstrated in Tab. 6, our code allocation scheme generally achieves the best overall performance, confirming the superior performance of our code allocation scheme.
@@ 7.2.3 Additional Ablation Study
Table 6: Ablation study on the code allocation scheme.
Effect of the model design. To verify the source of our performance improvement, we compare with a new “Baseline*", which has parameters comparable to our full model, achieved by increasing the ResBlock channel numbers of our image encoder and decoder. We present the results in Tab. 7. Our method significantly improves upon “Baseline*". The clear performance gap further confirms that the improvement comes from our model design rather than the increased parameters, indicating the effectiveness of our model design.
Method # Params (M) FID $\downarrow$ PSNR $\uparrow$ $\mathcal{L}_{1}$downarrow LPIPS $\downarrow$ AKD $\downarrow$ AED $\downarrow$
Baseline* 82.2 48.09 21.64 0.0549 0.2480 2.6798 0.2214
Ours 82.2 43.15 25.30 0.0355 0.1846 1.2039 0.1071
Table 7: Ablation study on the model design.
Codebook size. To assess how the codebook size affects the generation speed and quality, we vary the number of codes in the codebooks to achieve different codebook sizes and present results in Tab. 8. Larger codebooks generally improve generation quality by providing sufficient capacity to learn diverse motion and appearance codes, with only a slight decrease in speed/memory performance. A small codebook of 256 also performs well, likely because codes are retrieved more frequently during training, allowing for better optimization within the same training iterations. However, its image quality remains limited.
Number of Codes FID $\downarrow$ PSNR $\uparrow$ $\mathcal{L}_{1}$downarrow LPIPS $\downarrow$ AKD $\downarrow$ AED $\downarrow$ FLOPs (G) $\downarrow$ FPS $\uparrow$ Memory (M) $\downarrow$
256 47.50 25.18 0.0358 0.1861 1.1970 0.1039 351.57 15.60 6411
512 46.62 25.11 0.0362 0.1877 1.2190 0.1072 352.01 15.47 6411
1024 (Ours) 43.15 25.30 0.0355 0.1846 1.2039 0.1071 352.91 15.13 6413
Table 8: Ablation study on the codebook size. We present the results of different code numbers.
8 Limitation
A limitation of our method is the appearance leakage problem in cross-identity reenactment, where the face in the generated video tends to have a shape similar to that of the driving face rather than the source face. This issue arises from the keypoint-based motion flow estimator that we adopt to produce the initial coarse motion flow and the driving keypoints for multi-scale motion codebook compensation. Although this motion flow estimator is robust to non-facial motion, such as hair and neck movement, by learning unsupervised keypoints on talking heads, the keypoints also inherently model facial shapes, which leads to the entanglement of motion and shape. Thus, appearance leakage is a common issue for keypoint-based methods. Our method can effectively alleviate this issue by demonstrating better appearance preservation than other state-of-the-art keypoint-based approaches. As evidenced in Tab. 3, we achieve the highest CSIM score among these approaches, excluding LIA [31], which uses latent codes instead of keypoints for motion representation. This issue can also be mitigated through relative motion transfer [24], which is widely adopted by previous methods [16, 15, 25].